You invested money in two funds. Last year, the first fund paid a dividend of 8% and the second a dividend of 5%, and you received a total of $1330. This year, the first fund paid a 12% dividend and the second only 2% and you received a total of $1500. How much money did you invest in each fund?
Answer by mananth(16946)

Fund I x
Fund II ------------ y
0.08x+0.05y=1330 .............1
0.12x+0.02y=1500 .............2

multiply (1)by 5
Multiply (2) by 3
0.4 x + -0.02 y = 6650
0.36 x + 0.05 y = 4500

Add the two equations

0.76x=11150

/0.76

x=14671 Fund I

plug value of x in (1)

0.08x+0.05y=1330
1173.68 +0.05y= 1330
0.05y=1330-1173.68
0.05y= 156.32
y=3126 Fund II
neglect decimals